OpenStudy (mirandaregina):

(Multiple choice trig question) How do I evaluate cos(x)sin(x-90) + sin^2(x) ?

OpenStudy (mirandaregina):

A) 0, B) 1, C) 2x, D) 2sin(x)

Nnesha (nnesha):

use the identity \[\large\rm \sin(a+b)= \sin A *cosB - \cos A \sin B\] for sin(x-90)

OpenStudy (mirandaregina):

@Nnesha do I set it up as cos(x)sin(a + b) ?

Nnesha (nnesha):

\[\huge\rm cos(x)\color{ReD}{sin(x-90)} + \sin^2(x)\] just for red part use that formula i gave you sin(x-90) a=x and b = 90

Nnesha (nnesha):

i see i meant to say sin(a-b) not a+B

Nnesha (nnesha):

sorry about that let me retype it

Nnesha (nnesha):

use the identity \[\large\rm \sin(a-b)= \sin A *cosB - \cos A \sin B\] for sin(x-90) where a= x and b =90

Nnesha (nnesha):

so sin(x-90)=??

OpenStudy (mirandaregina):

@Nnesha sin(x-90) = -cos(x)

Nnesha (nnesha):

hmm right \[\huge\rm cos(x)*\color{ReD}{-cos(x)} + \sin^2(x)\] multiply cos(x) b -cos(x)

Nnesha (nnesha):

you're familiar with cos^2x + sin^2 x=1 identity solve this equation for sin^2(x)

OpenStudy (mirandaregina):

@Nnesha I have cos(x)(-cos(x)+sin^2 -cos^2(x)+sin^2(x) -cos(2x)

Nnesha (nnesha):

hmm 2nd line is correct use the identity to substitute sin^2(x) \[\rm \sin^2\theta+\cos^2 \theta =1\] solve this identity for sin^2(x)

Nnesha (nnesha):

\[\rm \sin^2\theta+\cos^2 \theta =1\] \[\sin^2 \theta = ???\]

OpenStudy (mirandaregina):

@Nnesha using the identity I have cos^2(x)+sin^2(x)+(-cos(x))cos(x) cos^2(x)+cos(x)(-cos(x)+sin^2(x) -cos^2(x)+cos^2(x)+sin^2(x) sin^2(x) + 0 sin^2(x)

Nnesha (nnesha):

hmm the first identity we used should be in form \[\sin(x-90)= \sin x * \cos 90 - \cos x *\sin 90 \]

Nnesha (nnesha):

how did you get -cos(x) and cos(x) at first line ?

Nnesha (nnesha):

were you talking about this identity ??? \[\rm \sin^2\theta+\cos^2 \theta =1\]

OpenStudy (mirandaregina):

@Nnesha yes

Nnesha (nnesha):

okay sin^2 x + cos^2 x =1 subtract cos^2 x both sides to solve for sin^2x

Nnesha (nnesha):

what do you get ? let m know ))

OpenStudy (mirandaregina):

@Nnesha sin^2 x + cos^2 x =1 sin^2x = -sin^2x -1 ?

Nnesha (nnesha):

 subtract both sides you will get \[\large\rm \sin^2x= 1-\cos^2x\]